Let W be the set of all 3×3 real diagonal matrices. Prove that W is a subspace of M33.

1
Expert's answer
2022-04-05T14:33:02-0400

Let "A,B \u2208 W" and "\u03b1\u2208\u211d" such that


"A=\begin{pmatrix}
 a_{11} & 0&0 \\
 0 & a_{22}&0\\0&0&a_{33}
\end{pmatrix}" "B=\begin{pmatrix}
 b_{11} & 0&0 \\
 0 & b_{22}&0\\0&0&b_{33}
\end{pmatrix}"


"A+αB=\begin{pmatrix}
 a_{11} & 0&0 \\
 0 & a_{22}&0\\0&0&a_{33}
\end{pmatrix} + α\begin{pmatrix}
 b_{11} & 0&0 \\
 0 & b_{22}&0\\0&0&b_{33}
\end{pmatrix}"



"=\begin{pmatrix}
 a_{11} & 0&0 \\
 0 & a_{22}&0\\0&0&a_{33}
\end{pmatrix} + \begin{pmatrix}
 αb_{11} & 0&0 \\
 0 & αb_{22}&0\\0&0&αb_{33}
\end{pmatrix}"



"=\begin{pmatrix}
 a_{11}+αb_{11} & 0&0 \\
 0 & a_{22}+αb_{22}&0\\0&0&a_{33}+αb_{33}
\end{pmatrix}"



Hence, "A+\u03b1B\u2208W"



Thus, "W" is a subspace of "M_{3\u00d73}(\u211d)"